Can someone help me on this problem?
I need to solve the following word problem and I need to show the equation you use for the solution.
A bus leaves a station at 1 p.m. traveling west at an average rate of 44 mi/h. One hour later a second bus leaves the same station, traveling east at a rate of 48 mi/h. At what time will the two buses by 274 mi apart?
Thanks, Ashley

Answer by venugopalramana(3286)

Can someone help me on this problem?
A bus leaves a station at 1 P.M., traveling west at an average rate of 44 mi/h. One hour later a second bus leaves the same station, traveling east at a rate of 48 mi/h. At what time will the two buses be 274 mi. apart?
Thanks,
Sher
RELATIVE SPEED OF 2 TRAINS GOING IN OPPOSITE DIRECTIONS = 44+48=92 MPH
INITIAL DISTANCE OF SEPERATION = DISTANCE TRAVELLED IN 1 HR BY THE FIRST BUS = 44 MILES
FINAL DISTANCE OF SEPERATION = 274 MILES
HENCE RELATIVE DISTANCE TRAVELLED AFTER SIMULTANEOUS START = 274-44=230
TIME NEEDED = 230/92 HRS = 2.5 HRS
HENCE AT 1+2.5+1=00 PM = 4.5 PM THE 2 BUSES WILL BE 274 MILES APART